OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4. 2 3 All real numbers No solution

OpenStudy (insa):

multiply them and then collect the like terms on each side and then find x its really easy and simple

OpenStudy (anonymous):

so it 3

OpenStudy (charlotte123):

u will get that the answer is all real numbers ^_^ to check, u can replace x with any number u can think of :D

OpenStudy (insa):

maybe i havent solved it

OpenStudy (charlotte123):

for proof: x: -2(x + 3) = -2(x + 1) - 4 (lets replace x with 1). so we get -2(1 + 3) = -2(1 + 1) - 4 --> -8 = -8. now lets substitute x with 2 :D we get -2(2 + 3) = -2(2 + 1) - 4 --> -10 = -10 ^_^ and so on :D so the answer is x = all real numbers :) @pinal145

OpenStudy (charlotte123):

@pinal145 got it? :D

OpenStudy (anonymous):

@charlotte123 yeh thanks

OpenStudy (charlotte123):

@pinal145 no prob :D welcome to OpenStudy :D if u feel like my answer is best, just click on that blue button for me :)

OpenStudy (charlotte123):

it has best response written on it :D u have a great day and enjoy ur stay here :) if u need any help just link me :D